On Mary's July water bill, the previous reading was 1237 ccf and the current reading was 1245 ccf. The water company charges $5.12 per ccf of water. In August, it was discovered that the July current reading was incorrect. It should have been 1241 ccf. How much of a credit will Mary receive for her over payment?

1
Expert's answer
2021-09-21T00:18:02-0400

incorrect current reading-previous reading =1245-1237=8ccf

water bill = "5.12\times 8=\$40.96"

calculation for correct water bill

correct current reading - previous reading=1241-1237=4ccf

correct water bill = "5.12\times 4=\$20.48"

credit received due to over payment =40.96-20.48="\$20.48"